Find the sum of the first 8 terms if a1=18 and d=3
Answer by jim_thompson5910(35256)

First we need the nth term

an = a1 + d(n-1)

an = 18 + 3(n-1)

an = 18 + 3n-3

an = 3n + 15

Now plug in n = 8 into the nth term formula to find the 8th term a8

an = 3n + 15

a8 = 3(8) + 15

a8 = 39

Now use the values for n, a1, an to find S in the formula below

S = n*(a1 + an)/2

S = 8*(a1 + a8)/2 ... there are 8 terms, so n = 8

S = 8*(18 + 39)/2

S = 8*(57)/2

S = 456/2

S = 228

So the sum of the first 8 terms is 228